OpenStudy (anonymous):

can anyone please help me Simplify the expression. the square root of the quantity 27 x to the fifth power, plus x times the square root of 3, plus nine times the square root of the quantity 3x

OpenStudy (anonymous):

can you? @e.cociuba

OpenStudy (anonymous):

can you? @e.cociuba \[\sqrt{27x^5} + x \sqrt{3} + 9\sqrt{3x}\]

OpenStudy (anonymous):

How would you like this "simplified"?

OpenStudy (anonymous):

do you need the answer choises? @elementalmagicks

OpenStudy (anonymous):

Anything helps.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no answer choices sorry @elementalmagicks